OpenStudy (anonymous):

Given the function f(x) = The quantity of 4x minus 2, divided by 3, which of the below expressions is correct? f−1(x) = The quantity of 2 minus 4x, divided by 3. f−1(x) = The quantity of 3x minus 2, divided by 4. f−1(x) = The quantity of negative 4x minus 2, divided by 3. f−1(x) = The quantity of 3x plus 2, divided by 4.

OpenStudy (anonymous):

@jim_thompson5910 @Josh21096

jimthompson5910 (jim_thompson5910):

First replace f(x) with y. Then swap x and y. Afterwards, solve for y. \[\Large f(x) = \frac{4x-2}{3}\] \[\Large y = \frac{4x-2}{3}\] \[\Large x = \frac{4y-2}{3}\] \[\Large 3x = 4y-2\] I'll let you finish up.

OpenStudy (anonymous):

im so confused dont u divide by 3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I multiplied both sides by 3 to move the 3 over. I ultimately want to solve for y.

OpenStudy (anonymous):

so subtract 4 from both sides

jimthompson5910 (jim_thompson5910):

why not add 2 to both sides 3x = 4y - 2 3x+2 = 4y 4y = 3x+2 y = ???

OpenStudy (anonymous):

y=3/4+2

OpenStudy (anonymous):

so it be D

jimthompson5910 (jim_thompson5910):

I think you meant to say \[\Large y = \frac{3x+2}{4}\]

OpenStudy (anonymous):

ya i just cant type it out like that

jimthompson5910 (jim_thompson5910):

alright, just checking, yes it is D

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome